A plane flies 720mi agains a steady 30-mi/h headwind and then returns to the same point with the wind. If the entire trip takes 10h, what is the plane's speed in still air?
Thank you so much

Found 2 solutions by Nate, checkley71:
Answer by Nate(3500)

If the plane keeps its own constant rate:
Rate against wind = planespeed - windspeed = 720
planespeed - 30 = 720
planespeed = 750mi

Answer by checkley71(8403)

720/(X+30)+720/(X-30)=10
[720(X-30)+720(X+30)]/(X+30)(X-30)=10
(720X-21600+720X+21600)/X^2-900=10
1440X/X^2-900=10
10X^2-9000=1440X
10X^2-1440X-9000=0
X^2-144X-900=0
(x-150)(x+6)=0
x-150=0
x=150 mph for the plane in still air.
proof
720/180+720/120)=10
4+6=10
10=10
###################################################
MERRY CHRISTMAS